# Finite Difference Solution Ansatz approach in Least-Squares Monte Carlo

### v7

@@ 2.2.5 Implementation Details
For dimension reduction at single-asset multi-factor level, the basic idea is to construct a 1D Markov process matching the marginal distribution of the original process, i.e., we take the conditional expectation of the instantaneous variance to construct the effective 1D local volatility:
@@
For Heston model without spot dependence on variance, we have
@@
Now we discuss the dimension reduction at multi-asset level. First of all, we can formally write down the effective 1D continuous dividend yield and local volatility in stochastic formats as
$q_{FD}(X,t)$ ::=:= $\operatorname{\mathbb{E}}\left[\frac{1}{X}\sum_{i}q_{i}\cdot\frac{\partial S_{B(W)}}{\partial S_{i}}\cdot S_{i}\bigg{|}\sum_{i}\frac{\partial S_{B(W)}}{\partial S_{i}}\cdot S_{i}=X\right],$
$\sigma^{2}_{FD}(X,t)$ ::=:= $\operatorname{\mathbb{E}}\left[\frac{1}{X^{2}}\sum_{ij}\rho_{ij}\sigma_{i}(S_{i},t)\sigma_{j}(S_{j},t)\frac{\partial S_{B(W)}}{\partial S_{i}}\frac{\partial S_{B(W)}}{\partial S_{j}}S_{i}S_{j}\bigg{|}\sum_{i}\frac{\partial S_{B(W)}}{\partial S_{i}}\cdot S_{i}=X\right],$
@@
For constant weighted basket with $X=S_{B}(t)$, there has been proposals to derive the effective local volatility approximately under the most general local volatility setting, see for example Avellaneda and Boyer-Olson, (2002). In the limit case of constant volatilities, i.e., Black-Scholes settings, we have below analytically tractable solution up to $2^{nd}$ order moment matching
@@
Plugging these into Eqn.2.16 as $q_{FD}=\bar{q}$ and $\sigma_{FD}=\bar{\sigma}$, we obtain the FD solution of continuation value $f^{FD,B}_{k}(x)$ after solving the PDE.
@@
    Follow the usual regression based procedures in LSM in Sec.2.1.2 to calculate the product PV using the basis functions (for each $t_{k}$) $\vec{\phi}_{R}^{\text{FD-LSM}}(x)$ concatenating $\{\hat{f}^{FD,B(W)}_{k}(x)\}$ and $\vec{\phi}_{R-1}^{\text{LSM}}(x)$.
@@ 3.1 Bermudan Option with $d=1$ under Black-Scholes
In Eqn.2.38, we use $r=3.96\%$ for numerical experiments in Bermudan options. Unless specified otherwise, the option maturity is set as $T=5$ with monthly exercise, i.e., $\Delta T=\frac{1}{12}$. In the simulation, we use $N_{R}=2^{13}$ and $N_{P}=2^{16}$ paths for regression and pricing stages, respectively, with weekly discretization steps. Sobol sequence is used in random number generation, which is shown as one of the most important variance reduction techniques (Areal and Armada, (2008)).
@@
As a further examination on the stability of the regression method, we show the $PV$ with arbitrary cutoff $R$ in Fig.3.2. Moving from low number of $R$ to medium level, e.g., $R=10$, we can see a significant improvement in terms of accuracy in LSM, where the $PV$ is more or less monotonically converging towards benchmark, largely and qualitatively consistent with previous literature (Frank J. Fabozzi, (2017)) focusing on put options, except the minor deterioration from $R=7$ to $R=8$ in the call option case; However, beyond $R=12$, the results deviate clearly away from benchmark in the call case, which is a clear sign of overfitting. On the other hand, for FD-LSM, the PV error against benchmark is bounded at 50 bps even up to $R=14$ for the call case. Although in this example the new regression method is exhibiting amazing robustness against overfitting, it is inadvisable to perform linear regression with high order of monomial terms in practice. From now on, we restrict and cap the cutoff to 5, unless specified otherwise.
@@ 3.2 Bermudan Option with $d=2$ and $d=4$ under Black-Scholes
Next, we further look at the case of $d=2$, where $\sigma_{1}=\sigma_{2}=30\%$ and $q_{1}=q_{2}=0$. For a fair comparison between FD-LSM and LSM, we use the same number of regressors with $R=4$ fixed: LSM makes use of full monomials up to the cubic term, whereas FD-LSM only uses perturbational monomials up to the quadratic term, apart form the main FD regressor. Looking at PV results as seen in Tab.3.2, the maximum of difference in FD-LSM against FD PDE2D benchmark  [^7] We use alternating direction implicit method (McKee and Mitchell, (1970)) in PDE2D. is merely 17 bps, whereas the counterpart in LSM is more than 2%, well beyond the simulation standard errors. Also, the computational overhead to achieve such accuracy is moderate as shown in the c.t. comparison. Last but not least, the Optimized Exercise Boundaries method (Opt-EB), as mentioned in Appendix A.4, can provide a valid independent check, although it tends to slightly underprice the option due to partial convergence in the external optimization.
@@ 3.3 Bermudan Option under Heston Stochastic Volatility
Here we consider Bermudan option pricing under Heston model as mentioned in Sec. 2.2.4. For Heston parameters, we use $r=0.02$, $q=0$, $\nu(0)=0.15$, $\kappa=5$, $\theta=0.16$, $\xi=0.9$ and $\rho^{S,\nu}=0.1$. We consider a monthly exercisable put option with $T=1$ and $K=1$; $N_{R}=2^{14}$ and $N_{P}=2^{18}$ are used in the simulation. In LSM, we use monomials regressors up to cubic terms, i.e., $\{S^{m}\nu^{n}|m+n\leq 3,n\in\mathbb{N}_{0},m\in\mathbb{N}_{0}\}$ and in FD-LSM, up to quadratic terms are used. We compare LSM and FD-LSM with the method “LSMC-PDE” proposed by Farahany et al., (2020) using the same settings. As we can see in Tab. 3.4, both FD-LSM and LSMC-PDE are returning PVs closer to PDE2D benchmark than classical LSM.
@@ 3.4 EPE/CVA for European option in $d=4$ under Black-Scholes
First, we inspect the EPE* profile, which is relevant to the CVA without considering WWR. In Fig.3.3, we plot the this profile using various cutoff $R$ under LSM and FD-LSM, where the s.e. is shown as error bar for indication. For LSM, EPE* deviates away from benchmark by up to $\sim$10 times of s.e. especially near the terminal end, with low level of $R=3$; by increasing $R$, the discrepancy is reduced in the longer end; however, at the same time, in the short end, there are spikes with abnormally sizable error bar arising around $t=1$ for $R=10$, indicting some kind of numerical instability.
@@
Now we continue to evaluate the total CVA using Eqn.2.15, taking into account WWR. As numerical settings, we use $R_{cpty}=0$; for the hazard rate, $a(t)=-4$ and $b=0.1$ are used, which are assumed to have been calibrated to counterparty credit spread using exact mark-to-market option values, and thus are independent of least-squares approximation. As seen in Tab. 3.5, as $R$ increases, CVA under LSM is exhibiting unstable move especially for $R\geq 10$, and the s.e. is showing explosive values up to 78 bps ($R=11$). By contrast, FD-LSM generates much more stable CVA all the way up to very high cutoff with a maximum s.e. capped at 27 bps ($R=11$).
@@ 3.5 WIC from $d=5$ to $d=50$ under Black-Scholes and Local Volatility
Finally, we also check the calculation time in last column of Tab. 3.7. In general, there is expected mild deterioration when switching to the new method, contributed by the additional PDE1D solvers and the cubic spline interpolation. However, the relative impacts are merely in single-digit range; thus we don’t expect this amount of impact would become a “showstopper” when applying the new regression scheme for generic option pricing and risk management from practitioners’ point of view.
@@
Local Volatility setting: We repeat previous numerical experiment under a local volatility setting with
@@ 4 Conclusion
In this manuscript, we presented a novel regression based approach for American-style option pricing. The formulism is to construct the ansatz to be regressed on using the 1D FD solution obtained from a backward PDE solver. Theoretical model-independent analysis indicates that the most effective least-square error reduction can be obtained by choosing a 1D payoff process with highest correlation against the original one. Effective dimension reduction methodologies to arrive at the 1D auxiliary process have also been discussed throughout various model settings, ranging from local volatility, to stochastic volatility (Heston model). Under this wide range of processes, numerical tests on Bermudan options and WIC indicate that FD-LSM produces accurate option prices benchmarking against FD-based full PDE approach in low dimensions ($d\leq 2$), and consistently overshadows classical LSM under high dimensions (up to 50). By stressing the market and contractual parameters, we showed that the combination of the ad-hoc FD-based ansatz and perturbational monomials as regressors, gives rise to stable results especially with exercise frontiers far away from the ATM region. At the same time, there has not been noticeable degradation in terms of the computational efficiency with the additional calculation in PDE solver and spline interpolation. This effective and efficient method can be implemented as a generic framework for pricing and risk management on structure derivative products with American-style exercise features. Going beyond option pricing, since LSM is also widely used to approximate the future exposure for products that cannot be valued analytically in the context of XVA, the new method is applicable there for the sake of accuracy and numerical stability.
@@ A.3 Regression Approach in Monte Carlo framework
A.3.1 Least-Squares Solution
Translating objective function to be minimized in Eqn.2.3 into the language within Monte Carlo framework, we have
@@ A.3.1 Least-Squares Solution
After some algebra, the solution of the minimization problem can be written as:
$\vec{\beta}^{P}_{k}=(\mathbb{X^{P}}^{T}_{k}\mathbb{X^{P}}_{k})^{-1}\mathbb{X^{P}}^{T}_{k}\vec{Y},$ (4.1)
@@
A.3.2 Local Regression Error
For the sake of further analysis on the regression error due to the approximation, we re-write the exact conditional expectation $\vec{Y}^{0}_{k}$ as
$\vec{Y}^{0}_{k}=\mathbb{X^{P}}_{k}\vec{\beta}^{0}_{k}$
with the exact projection coefficients denoted by $\vec{\beta}^{0}_{k}$, as opposed to $\vec{\beta}^{P}_{k}$ under Eqn.2.2.
Then squared difference between $P_{k}[Y]$ and $Y^{0}_{k}$ can be written, in the linear algebra representation following the notations introduced in Eqn.4.1, as
$\begin{split}\operatorname{\mathbb{E}}[(P_{k}[Y]-Y^{0}_{k})^{2}|\operatorname{\mathcal{F}}_{0}]&=[\mathbb{X^{P}}_{k}\cdot\Delta\vec{\beta}^{P}_{k}]^{T}[\mathbb{X^{P}}_{k}\cdot\Delta\vec{\beta}^{P}_{k}],\\ \end{split}$ (4.2)
where
$\begin{split}\Delta\vec{\beta}^{P}_{k}:&=\vec{\beta}^{P}_{k}-\vec{\beta}^{0}_{k}\\ &=(\mathbb{X^{P}}^{T}_{k}\mathbb{X^{P}}_{k})^{-1}\mathbb{X^{P}}_{k}^{T}(\vec{Y}-\vec{Y}^{0}_{k}).\end{split}$ (4.3)
Expanding Eqn.4.2 with the help of basic linear algebra properties that are also used by Newey, (1997) in the their proof of Theorem 1, we have
$\begin{split}\operatorname{\mathbb{E}}[(P_{k}[Y]-Y^{0}_{k})^{2}|\operatorname{\mathcal{F}}_{0}]&=(\vec{Y}-\vec{Y}^{0}_{k})^{T}\mathbb{X^{P}}_{k}(\mathbb{X^{P}}^{T}_{k}\mathbb{X^{P}}_{k})^{-1}\mathbb{X^{P}}^{T}_{k}\mathbb{X^{P}}_{k}(\mathbb{X^{P}}^{T}_{k}\mathbb{X^{P}}_{k})^{-1}\mathbb{X^{P}}^{T}_{k}(\vec{Y}-\vec{Y}^{0}_{k}),\\ &=\text{tr}(\mathbb{X^{P}}_{k}(\mathbb{X^{P}}^{T}_{k}\mathbb{X^{P}}_{k})^{-1}\mathbb{X^{P}}^{T}_{k}[(\vec{Y}-\vec{Y}^{0}_{k})\otimes(\vec{Y}-\vec{Y}^{0}_{k})^{T}]),\\ &\leq c^{P}_{k}\cdot\text{tr}[(\vec{Y}-\vec{Y}^{0}_{k})\otimes(\vec{Y}-\vec{Y}^{0}_{k})^{T}],\\ &=c^{P}_{k}\cdot\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(Y|\operatorname{\mathcal{F}}_{t_{k}})|\operatorname{\mathcal{F}}_{0}],\end{split}$ (4.4)
where the submultiplicative property of matrix trace (tr), implied by the Cauchy-Schwarz inequality, has been applied in the derivation and $c^{P}_{k}$ is a constant depending on the projection method and its details (e.g. number of cutoffs $R$ and paths $N_{R}$). Eqn.4.4 relates the upper bound of the local regression error on $Y$ to its variance.
A.4 Optimized Exercise Boundaries (Opt-EB)

### v8

@@ 2.3.4 Implementation Details
At single-asset multi-factor level, the basic idea is to construct a 1D Markov process matching the marginal distribution of the original process, i.e., we take the conditional expectation of the instantaneous variance to construct the effective 1D local volatility:
@@
For Heston model in Eqn.2.2 without spot dependence on variance, we have
@@
Now we discuss the dimension reduction at multi-asset level. First of all, we can formally write down the effective 1D continuous dividend yield and local volatility as (see Avellaneda et al., (2002))
$\begin{split}q_{FD}(X,t)&:=\operatorname{\mathbb{E}}\left[\frac{1}{X}\sum_{i}q_{i}\cdot\frac{\partial S_{B(W)}}{\partial S_{i}}\cdot S_{i}\bigg{|}\sum_{i}\frac{\partial S_{B(W)}}{\partial S_{i}}\cdot S_{i}=X\right],\\ \sigma^{2}_{FD}(X,t)&:=\operatorname{\mathbb{E}}\left[\frac{1}{X^{2}}\sum_{ij}\rho_{ij}\sigma_{i}(S_{i},t)\sigma_{j}(S_{j},t)\frac{\partial S_{B(W)}}{\partial S_{i}}\frac{\partial S_{B(W)}}{\partial S_{j}}S_{i}S_{j}\bigg{|}\sum_{i}\frac{\partial S_{B(W)}}{\partial S_{i}}\cdot S_{i}=X\right],\end{split}$ (2.60)
@@
For constant weighted basket with $X=S_{B}(t)$, there has been proposals to derive the effective local volatility approximately under the most general local volatility setting, see for example Avellaneda et al., (2002). In the limit case of constant volatilities, i.e., Black-Scholes settings, we have below analytically tractable solution up to $2^{nd}$ order moment matching
@@
Plugging these into Eqn.2.20 as $q_{FD}=\bar{q}$ and $\sigma_{FD}=\bar{\sigma}$, we obtain the FD solution of continuation value $f^{FD,B}_{k}(x)$ after solving the PDE.
@@
    Follow the usual regression based procedures in LSM in Sec.2.2.2 to calculate the product PV using the basis functions (for each $t_{k}$) $\vec{\phi}_{R}^{P^{\text{FD-LSM}}}(x)$ concatenating $\{\hat{f}^{FD,B(W)}_{k}(x)\}$ and $\vec{\phi}_{R-1}^{P^{\text{LSM}}}(x)$.
@@ 3.1 Bermudan Option with $d=1$ under Black-Scholes
In Eqn.2.1, we use $r=3.96\%$ for numerical experiments in Bermudan options. Unless specified otherwise, the option maturity is set as $T=5$ with monthly exercise, i.e., $\Delta T=\frac{1}{12}$. In the simulation, we use $N_{R}=2^{13}$ and $N_{P}=2^{16}$ paths for regression and pricing stages, respectively, with weekly discretization steps. Sobol sequence is used in random number generation, which is shown as one of the most important variance reduction techniques (Areal and Armada, (2008)).
@@
As a further examination on the stability of the regression method, we show the $PV$ with arbitrary cutoff $R$ in Fig.3.2. Moving from low number of $R$ to medium level, e.g., $R=10$, we can see a significant improvement in terms of accuracy in LSM, where the $PV$ is more or less monotonically converging towards benchmark, largely and qualitatively consistent with previous literature (Frank J. Fabozzi, (2017)) focusing on put options, except the minor deterioration from $R=7$ to $R=8$ in the call option case; However, beyond $R=12$, the results deviate clearly away from benchmark in the call case, which is a clear sign of overfitting. On the other hand, for FD-LSM, the PV error against benchmark is bounded at 50 bps even up to $R=14$ for the call case. Although in this example the new regression method is exhibiting amazing robustness against overfitting, it is inadvisable to perform linear regression with high order of monomial terms in practice. From now on, we restrict and cap the cutoff to 4, unless specified otherwise.
@@ 3.2 Bermudan Option with $d=2$ and $d=4$ under Black-Scholes
Next, we further look at the case of $d=2$, where $\sigma_{1}=\sigma_{2}=30\%$ and $q_{1}=q_{2}=0$. For a fair comparison between FD-LSM and LSM, we use the same number of regressors with $R=4$ fixed: LSM makes use of full monomials up to the cubic term, whereas FD-LSM only uses perturbational monomials up to the quadratic term, apart form the main FD regressor. Looking at PV results as seen in Tab.3.2, the maximum of difference in FD-LSM against FD PDE2D benchmark  [^8] We use alternating direction implicit method (McKee and Mitchell, (1970)) in PDE2D. is merely 17 bps, whereas the counterpart in LSM is more than 2%, well beyond the simulation standard errors. Also, the computational overhead (single-digit in percentage) to achieve such accuracy is moderate as shown in the c.t. comparison against LSM. The PDE2D solver’s c.t, approximately 25% of that required by LSM, is non-negligible compared to the 1D calculations. This also justifies our choice of using the 1D PDE as the ansatz. Last but not least, the Optimized Exercise Boundaries method (Opt-EB), as mentioned in Appendix A.5, can provide a valid independent check, although it tends to slightly underprice the option due to partial convergence in the external optimization.
@@ 3.3 Bermudan Option under Heston Stochastic Volatility
Here we consider Bermudan option pricing under Heston model as mentioned in Sec. 2.1. For Heston parameters, we use $r=0.02$, $q=0$, $\nu(0)=0.15$, $\kappa=5$, $\theta=0.16$, $\xi=0.9$ and $\rho^{S,\nu}=0.1$. We consider a monthly exercisable put option with $T=1$ and $K=1$; $N_{R}=2^{14}$ and $N_{P}=2^{18}$ are used in the simulation. In LSM, we use monomials regressors up to cubic terms, i.e., $\{S^{m}\nu^{n}|m+n\leq 3,n\in\mathbb{N}_{0},m\in\mathbb{N}_{0}\}$ and in FD-LSM, up to quadratic terms are used. We compare LSM and FD-LSM with the method “LSMC-PDE” proposed by Farahany et al., (2020) using the same settings. As we can see in Tab. 3.4, both FD-LSM and LSMC-PDE are returning PVs closer to PDE2D benchmark than classical LSM.
@@ 3.4 EPE/CVA for European option in $d=4$ under Black-Scholes
First, we inspect the EPE* profile, which is relevant to the CVA without considering Wrong Way Risk (WWR). In Fig.3.3, we plot the this profile using various cutoff $R$ under LSM and FD-LSM, where the s.e. is shown as error bar for indication. For LSM, EPE* deviates away from benchmark by up to $\sim$10 times of s.e. especially near the terminal end, with low level of $R=3$; by increasing $R$, the discrepancy is reduced in the longer end; however, at the same time, in the short end, there are spikes with abnormally sizable error bar arising around $t=1$ for $R=10$, indicting some kind of numerical instability.
@@
Now we continue to evaluate the total CVA using Eqn.2.19.
To further take into account the WWR, we consider a hazard rate with below functional dependence on the option mark-to-market value $\tilde{F}(\omega,t)$(Hull and White, (2012)):
$h_{cpty}(\tilde{F}(\omega,t),t)=\ln[1+e^{a_{WWR}(t)+b_{WWR}\tilde{F}(\omega,t)}],$
where $a_{WWR}(t)$ and $b_{WWR}$ are WWR model parameters.
As numerical settings, we use $R_{cpty}=0$; for the hazard rate, $a_{WWR}(t)=-4$ and $b_{WWR}=0.1$ are used, which are assumed to have been calibrated to counterparty credit spread using exact mark-to-market option values, and thus are independent of least-squares approximation. As seen in Tab. 3.5, as $R$ increases, CVA under LSM is exhibiting unstable move especially for $R\geq 10$, and the s.e. is showing explosive values up to 78 bps ($R=11$). By contrast, FD-LSM generates much more stable CVA all the way up to very high cutoff with a maximum s.e. capped at 27 bps ($R=11$).
@@ 3.5 WIC from $d=5$ to $d=50$ under Black-Scholes and Local Volatility
Finally, we also check the calculation time in last column of Tab. 3.7. In general, there is expected mild deterioration when switching to the new method, contributed by the additional PDE1D solvers and the cubic spline interpolation. However, the relative impacts are merely in single-digit range; thus we don’t expect this amount of impact would become a “showstopper” when applying the new regression scheme for generic option pricing and risk management from practitioners’ point of view. Local Volatility setting: We repeat previous numerical experiment under a local volatility setting with
@@ 4 Conclusion
In this manuscript, we presented a novel regression based approach for American-style option pricing. The formulism is to construct the ansatz to be regressed on using the 1D FD solution obtained from a backward PDE solver. Theoretical analysis indicates that the most effective error reduction can be obtained by choosing a 1D payoff process with highest correlation against the original one. Effective dimension reduction methodologies to arrive at the 1D auxiliary process have also been discussed throughout various model settings, ranging from local volatility to stochastic volatility (Heston model). Under this wide range of processes, numerical tests on Bermudan options and WIC indicate that FD-LSM produces accurate option prices benchmarking against FD-based full PDE approach in low dimensions ($d\leq 2$), and consistently overshadows classical LSM under high dimensions (up to 50). By stressing the market and contractual parameters, we showed that the combination of the ad-hoc FD-based ansatz and perturbational monomials as regressors, gives rise to stable results especially with exercise frontiers far away from the ATM region. At the same time, there has not been noticeable degradation in terms of the computational efficiency with the additional calculation in PDE solver and spline interpolation; the increase in runtime is primarily single-digit percentage, and only reaches low double digits in exceptional cases, see e.g. the last column of Tab.3.8. This effective and efficient method can be implemented as a generic framework for pricing and risk management on structure derivative products with American-style exercise features. Going beyond option pricing, since LSM is also widely used to approximate the future exposure for products that cannot be valued analytically in the context of XVA, the new method is applicable there for the sake of accuracy and numerical stability.
@@ A.3 Regression Approach in Monte Carlo framework
Translating objective function to be minimized in Eqn.2.5 into the language within Monte Carlo framework, we have
@@
After some algebra, the Monte Carlo estimator for the solution of the minimization problem can be written as:
$\hat{\vec{\beta}}^{P}_{k}=(\mathbb{X^{P}}^{T}_{k}\mathbb{X^{P}}_{k})^{-1}\mathbb{X^{P}}^{T}_{k}\vec{Y},$ (4.1)
@@
A.4 Local Regression Error in LSM
To begin with, we analyze the MSE for LSM by writing down the exact projection coefficients as
$\vec{\beta}^{P^{\text{LSM}}0}_{k}=\vec{\beta}^{P^{\text{LSM}}}_{k}(F_{k})=(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1}\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}F_{k}].$ (4.2)
Introducing $\Delta\vec{\beta}^{P^{\text{LSM}}}_{k}=\vec{\beta}^{P^{\text{LSM}}}_{k}(e^{-r\Delta T}V_{k+1})-\vec{\beta}^{P^{\text{LSM}}0}_{k}=\vec{\beta}^{P^{\text{LSM}}}_{k}(e^{-r\Delta T}V_{k+1}-F_{k})$, where $\vec{\beta}^{P^{\text{LSM}}}_{k}(\cdot)$ is obtained from Eqn.2.8 after re-labeling $P$ to $P^{\text{LSM}}$, the local error is derived as
$\begin{split}&\operatorname{\mathbb{E}}[(e^{-r\Delta T}V_{k+1}-F_{k})^{2}]\\ =&\operatorname{\mathbb{E}}[(\Delta\vec{\beta}^{P^{\text{LSM}}}_{k}\cdot\vec{\phi}^{P^{\text{LSM}}}_{R,k})^{2}]\\ =&\operatorname{\mathbb{E}}[\text{tr}(\Delta\vec{\beta}^{P^{\text{LSM}}T}_{k}\vec{\phi}^{P^{\text{LSM}}}_{R,k}\Delta\vec{\beta}^{P^{\text{LSM}}T}_{k}\vec{\phi}^{P^{\text{LSM}}}_{R,k})]\\ =&\operatorname{\mathbb{E}}[\text{tr}(\vec{\phi}^{P^{\text{LSM}}T}_{R,k}(e^{-r\Delta T}V_{k+1}-F_{k})^{2}(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1}\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}\\ \quad&(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1}\vec{\phi}^{P^{\text{LSM}}}_{R,k})]\qquad\qquad\qquad\qquad\quad\text{ (Law of Iterated Expectation)}\\ =&\operatorname{\mathbb{E}}[(e^{-r\Delta T}V_{k+1}-F_{k})^{2}\text{tr}(\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1}\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}\\ \quad&(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1})]\qquad\qquad\qquad\qquad\qquad\qquad\text{(Cyclic property of matrix trace)}\\ \leq&\operatorname{\mathbb{E}}[(e^{-r\Delta T}V_{k+1}-F_{k})^{2}]\cdot\operatorname{\mathbb{E}}[\text{tr}(\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1}\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}\\ \quad&(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1})]\qquad\qquad\qquad\qquad\qquad\qquad\text{(Cauchy-Schwarz inequality)}\\ \leq&\operatorname{\mathbb{E}}[(e^{-r\Delta T}V_{k+1}-F_{k})^{2}]\cdot\text{tr}(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}](\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1}\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}]\\ \quad&(\operatorname{\mathbb{E}}[\vec{\phi}^{P^{\text{LSM}}}_{R,k}\vec{\phi}^{P^{\text{LSM}}T}_{R,k}])^{-1})\qquad\qquad\qquad\qquad\qquad\qquad\text{ (Cauchy-Schwarz inequality)}\\ =&\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})]\cdot\text{tr}({\bf{I}}_{R})\\ =&\operatorname{\mathbb{E}}[\operatorname{\mathrm{Var}}(e^{-r\Delta T}V_{k+1}|\operatorname{\mathcal{F}}_{t_{k}})]\cdot R,\end{split}$ (4.3)
with ${\bf{I}}_{R}$ being $R\times R$ identity matrix. While the above derivation is similar to the proof of Theorem 1 in Newey, (1997), we would emphasize specifically that the tightness of the first occurrence when applying Cauchy-Schwarz inequality is justified by $\vec{\phi}^{P^{\text{LSM}}}_{R,k}$ and $V_{k+1}$ being agnostic to each other.
A.5 Optimized Exercise Boundaries (Opt-EB)